# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v1

@@ 5.4 Thermalization Ablation Study
We ablate our different regularizers in Eq. (1)-(3) using the Fréchet Inception Distance (FID) [19]. The FID measures image distribution similarity between a real and a synthetic set of images based on an Inception v3 [51] model pre-trained on ImageNet [10]. For our evaluation, we report the average FID between the complete SEJONG dataset and twenty equal-sized samples from the FAKE dataset. Here, we use the SEJONG dataset as a reference thermal dataset to evaluate the similarity of differently thermalized images to true thermal images. We thermalize each sample with i) our model without regularization, ii) our model with patch-based regularization (Eq. 2) only, iii) our model with segmentation-based regularization (Eq. 3) only, and iv) our final model. All models are trained with the same setup and we randomly choose the ‘warm’ or the ‘cold’ variant for each image. Table 4 shows the average FIDs and the standard deviations. We see that the segmentation-based regularizer leads to a large boost in perceptual similarity and that the patch-based regularizer leads to a smaller boost. The combination of both regularizers gives the best result and confirms the effectiveness of our RGB2Thermal loss.
5.5 Landmarking Ablation Study
To study the impact of our thermal data, we additionally report the CHARLOTTE results of our landmarker trained with i) RGB images only and ii) thermal images only. Again, we use label adaptation for all variations. Table 5 shows the results. Training with the T-FAKE dataset significantly improves the accuracy of thermal landmarking across all conditions. In addition, multimodal training with the FAKE and T-FAKE datasets leads to better thermal landmarking performance than training only with the T-FAKE dataset. Moreover, we ablate our RGB performance on 300W by testing the model trained on RGB images only, see Table 3. Here, we see a slight drop in performance caused by multimodal training. Nevertheless, our multimodal model still outperforms FA [6] and 3FabRec [5].
@@ A.2 Thermalization Implementation Details
We use a U-Net $T_{\theta}$ with a Resnet34 encoder pre-trained on ImageNet [10] and train it for 10 epochs with SEJONG batches of size $64$ and FAKE batches of size $64$. Further, we use an Adam optimizer with an initial learning rate of 0.001 which we reduce to 0.0001 after 4 epochs. Based on a random split, we use 80% of the SEJONG data for training and all available FAKE data. Also, we use the geomloss [11] implementation for the Wasserstein patch loss with $\lambda_{E}=0.1^{6}$. We normalize the squared error loss $\|\cdot-\cdot\|_{2}^{2}$ by dividing with the image dimension, here $256^{2}$. Based on the $5$ scales and patch dimension $8^{2}$, we set $\lambda_{W}=0.01C$ for our final model with the normalization constant $C=(5\cdot 8^{2})^{-1}$. We set $\lambda_{R}=1$. Note that the choice of $\lambda_{R}=1$ and $C$ are motivated to control the value range. The (normalized) MSE data fidelity term $\|\cdot-\cdot\|_{2}^{2}$ which we evaluate on normalized SEJONG images takes values in the range $[0,1]$ for arbitrary images with pixel range $[0,1]$ due to dimensional normalization. To have a similar value range for the evaluated FAKE images, the patch-based regularizer $W$ takes values in $[0,1]$ for $\lambda_{W}=C$ with arbitrary normalized images. Moreover, the segmentation-based regularizer $R$ also takes values in $[0,1]$ for such images given arbitrary normalized reference temperature values in $[0,1]$ for $\lambda_{R}=1$.
@@ A.3 SEJONG Dataset
The SEJONG dataset illustrates the impact of various disguises, including glasses, wigs, and fake beards. Each subject in the dataset is presented with different disguises. As a result, it includes a lot of clothing and hairstyle variations. Most participants have a Southeast Asian or Central Asian ethnic background, but people of other ethnicities are included too. The number of participants identifying as male or female is balanced. This makes it an attractive dataset candidate for thermalization training that is supposed to generalize to a large variety of subjects. Due to reasons of data privacy, we refer to the original publication [7] and abstain from showing SEJONG images.
A.4 Extended Thermalization Ablation Study
@@
Given that the regularizers are solely defined for the synthetic images, we fix $\lambda_{T}=1$ to ensure that the regularizer is on the same scale as the MSE of the real images and vary $\lambda_{W}$ for our ablation. We train all models with the same setup and random seeds. The FID implementation is the default PyTorch-Ignite [15] implementation. Here, we extend the ablation study table in the main document which only displays the best result for $\lambda_{W}=0.01C$. In particular, we display the results of a grid search for $\lambda_{W}$ and $\lambda_{R}$ in Table 7. As described in the main document, we calculate the mean FID and its standard deviation for twenty different subsets of the T-FAKE dataset. Moreover, we use the same setup to compare the perceptual quality of the ‘cold’ and the ‘warm’ setup of our final T-FAKE dataset, see Table 8. Figure 5 shows T-FAKE samples with the ‘cold’ and ‘warm’ variants. In addition, we present some samples generated with different regularization configurations in Figure 6 for the ‘cold’ setup and in Figure 7 for the ‘warm’ setup. Here, we can visually see the impact of the different regularizers.
@@
$\lambda_{W}=1C$ 0.2714 $\pm$ 0.0045 0.3139 $\pm$ 0.0035
$\lambda_{W}=0.1C$ 0.2710 $\pm$ 0.0042 0.3134 $\pm$ 0.0048
$\lambda_{W}=0.01C$ 0.2666 $\pm$ 0.0037 0.3120 $\pm$ 0.0043
$\lambda_{W}=0$ 0.2719 $\pm$ 0.0054 0.3325 $\pm$ 0.0039
Table 7: Impact of the regularization parameters on the perceptual quality measured with the FID with $C=(5\cdot 8^{2})^{-1}$.
@@
Figure 6: Regularization impact on images for ‘cold’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.001C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
Figure 7: Regularization impact on images for ‘warm’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.001C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
@@ B.2 Landmarking Implementation Details
For training the landmarker $T_{\psi}$, we finetune a model that has been pre-trained on the original FAKE dataset for the sparse landmarker. For the dense landmarker, pre-training happened on the original FAKE dataset including a first version of the T-FAKE dataset as augmentation with a probability of 0.2. During refinement, thermal images are used with a probability of 0.4 split with equal probability for ‘cold’ and ‘warm’ conditions. We finetune for 100 epochs with a learning rate of 0.0004, Adam optimizer with weight decay, batch size of 512, and OneCycleLR scheduler.
@@ B.3 Label Adaptation Implementation Details
Label Adaptation was trained for each method on the predictions on all detected faces on a random 1000-image CHARLOTTE split. We train a model $T_{\zeta}$ for 2000 epochs with a learning rate of 0.002, OneCycleLR, and Adam optimizer. As landmark augmentations, we apply random rotation up to 45° as well as random shearing during training. The label adaptation network is a five-layer perceptron with fully connected layers that takes the predicted landmarks together with the resize factor as input. The latter accounts for varying degrees of quantization at different image sizes in the CHARLOTTE ground truth. The label adaptation generally handles even outlier predictions but can also contain fail cases, (see Figure 10).
@@
RGB Model Inference. We use two different pre-processing approaches to include landmarkers solely developed for RGB images in the evaluation. Firstly, gray-value images, where the temperature between 20° and 45° is normalized and, secondly the pre-processing stack proposed in [12]. The pre-processing stack consists of temperature clamping between 20°C and 45°C, unsharp masking with two sets of parameters with and without temperature inversion. The reported landmarks are the averages over all detected faces. This simple pre-processing stack is a simple method for boosting RGB landmarker performance for thermal images [12]. As a result, we can include RGB landmarkers as a baseline for thermal landmarking models.
B.4 Inference Ablation
@@
We analyze the impact of our inference strategy, see Table 9. We compare inference on i.) the complete image scaled to $224\times 224$ (whole image), ii.) followed by refinement on a bounding box computed from the predictions obtained with i.) and finally iii.) with the sliding window approach described in the main document. Method iii.) produces the best results over all images except for Charlotte low, where ii.) performs slightly better while at the same time being also suitable for real-time estimation.
@@
Figure 9: Results on examples from the CHARLOTTE dataset [3] with different RGB and thermal predictors and our models. For images without landmarks, no faces were detected. The performance of RGB methods can be greatly improved when the images are inverted or sharpened indicated by *. The first column shows the limitations in the CHARLOTTE ground truth: profile annotation convention for frontal views (1st row), quantization artifacts for low-resolution images (4th row), and translated annotations (last row).
Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
Sliding window ($\bar{\sigma}<\infty$) 0.0740 0.1346 0.0684 0.1420 0.1051
Refined bounding box ($\bar{\sigma}<\infty$) 0.0847 0.1329 0.0696 0.1494 0.1095
Whole image ($\bar{\sigma}<\infty$) 0.1091 0.1868 0.0842 0.2139 0.1490
@@
Table 9: Different strategies for landmark computation. The final results are estimated with sliding windows similar to [53], however, we achieved comparable results when computing the landmarks on input images rescaled to $224\times 224$. Here, we do not exclude high-uncertainty images and evaluate all images without failure rate, i.e., $\bar{\sigma}<\infty$.
@@ Appendix C Large-Scale Visualization
For a large number of T-FAKE samples, we refer to Figures 11 and 12. Here, we simply use the first 128 images based on the numerical naming convention of the original FAKE dataset.

### v2

@@ 5.5 Thermalization Ablation Study
As an additional ablation of our regularizers, Table 4 presents all the previously introduced metrics for different choices of $\lambda_{W}$ and $\lambda_{T}$. We see that the segmentation-based regularizer leads to a large boost in perceptual similarity and temperature prediction power. The patch-based regularizer leads to a smaller boost. The combination of both regularizers always gives the best or second-best result. This confirms the effectiveness of our RGB2Thermal loss.
@@ A.2 Thermalization Implementation Details
For the patch-based regularizers, we use a patch size of $8$. We disregard the background by excluding synthetic patches based on the ground truth background segmentation and completely black real patches. For our multi-scale approach, we sum the regularizer over $5$ scales with a downsampling factor of $0.5$. Furthermore, facial temperatures depend on the surrounding temperature. Thus, we train a ‘warm’ and a ‘cold’ model with different reference temperature values for the segmentation-based regularizers. For data augmentation, we use the same random rotations and cropping for the natural RGB and thermal images. Moreover, we apply random color changes, blurring, and shadow augmentations [43] exclusively to the natural RGB images. Next, we apply random rotations and cropping for the synthetic RGB images. Here, we also fill holes in the original ‘glasses’ segmentation masks showing outlines of the frame only to highlight transparent, but heat-blocking glass or plastic. Lastly, we replace the original background with a black background based on the known semantic segmentation. We use a U-Net $T_{\theta}$ with a Resnet34 encoder pre-trained on ImageNet [10] and train it for 10 FAKE epochs with SEJONG batches of size $64$ and FAKE batches of size $64$. This corresponds to approximately 100 SEJONG epochs. Further, we use an Adam optimizer with an initial learning rate of 0.001 which we reduce to 0.0001 after 4 epochs. Based on a random split, we use 80% of the SEJONG data for training and all available FAKE data. Also, we use the geomloss [12] implementation for the Wasserstein patch loss with $\lambda_{E}=0.1^{6}$. We normalize the squared error loss $\|\cdot-\cdot\|_{2}^{2}$ by dividing with the image dimension, here $256^{2}$. Based on the $5$ scales and patch dimension $8^{2}$, we set $\lambda_{W}=0.01C$ for our final model with the normalization constant $C=(5\cdot 8^{2})^{-1}$. We set $\lambda_{R}=1$. Note that the choice of $\lambda_{R}=1$ and $C$ are motivated to control the value range. The (normalized) MSE data fidelity term $\|\cdot-\cdot\|_{2}^{2}$ which we evaluate on normalized SEJONG images takes values in the range $[0,1]$ for arbitrary images with pixel range $[0,1]$ due to dimensional normalization. To have a similar value range for the evaluated FAKE images, the patch-based regularizer $W$ takes values in $[0,1]$ for $\lambda_{W}=C$ with arbitrary normalized images. Moreover, the segmentation-based regularizer $R$ also takes values in $[0,1]$ for such images given arbitrary normalized reference temperature values in $[0,1]$ for $\lambda_{R}=1$.
@@ A.3 SEJONG Dataset
The SEJONG dataset illustrates the impact of various disguises, including glasses, wigs, and fake beards. Each subject in the dataset is presented with different disguises. As a result, it includes a lot of clothing and hairstyle variations. Most participants have a Southeast Asian or Central Asian ethnic background, but people of other ethnicities are included too. The number of participants identifying as male or female is balanced. This makes it an attractive dataset candidate for thermalization training that is supposed to generalize to a large variety of subjects. Due to reasons of data privacy, we refer to the original publication [7] and have to abstain from showing randomly selected SEJONG images.
A.4 Generalization to Out-of-Lab Conditions
The success of neural networks in the last two decades has been tremendous, especially in the imaging domain. Nevertheless, a common empirical finding is the limited capability of neural networks to generalize to new settings [47]. Often, this limitation is caused by biased datasets. Particularly in the biomedical domain, data acquisition is a valuable task. However, due to real-world restrictions, data is often acquired in laboratory conditions. Acquiring paired RGB and thermal facial images requires a calibrated multimodal camera setup. Most multimodal facial datasets are restricted to frontal views with relatively neural expressions and frontal lighting, e.g. [48, 7, 13]. As a direct consequence, most Thermal2RGB research has focused on learning the transformation purely for frontal images with frontal lighting at room temperature [56, 64, 45]. To our knowledge, our regularized model is the first facial model to promote the explicit generalization of a learned thermal transformation to new poses, facial expressions, and lighting conditions and to simulate distinct temperature conditions. However, due to a lack of paired multimodal ‘in-the-wild’ facial datasets, we have to partially resort to metrics for unsupervised image translation, i.e., the FID [21]. To additionally visualize this result, we show results for the TUFTS [48] dataset containing RGB and thermal images. Again, we find that paired RGB and images are only available for frontal views. We present the results of applying our baseline model without regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), a Pix2Pix model, and our final ‘cold’ model to the dataset subset with paired RGB images in Figure 6. We chose the ‘cold’ model because the images were recorded at room temperature. We see only a marginal impact of our regularization on the prediction. However, the TUFTS dataset additionally contains RGB images without paired thermal images recorded from different angles. Therefore, we also apply both models to random images taken from a fixed side angle. The result is displayed in Figure 7. We see that the unregularized model generates various large facial artifacts whereas our final model contains almost no facial artifacts. This shows that the main limitation of training RGB2Thermal and Thermal2RGB models is the lack of paired ‘in-the-wild’ multimodal images. Our regularization allows us to overcome this limitation.
Figure 6: Thermalization results for frontal RGB images with paired thermal images from the TUFTS [48] database without regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), Pix2Pix and our final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
Figure 7: Thermalization results for side RGB images without paired thermal images from the TUFTS [48] database without regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), Pix2Pix and our final model ($\lambda_{W}=0.001C$, $\lambda_{R}=1$) (top to bottom).
A.5 Thermalization Comparison Details
[…]
Figure 8: Comparison of FAKE images thermalized with Pix2Pix [26], CycleGAN [66], CUT [49], QS-Attn [23], our supervised baseline and our final model (top to bottom, background removed).
Figure 9: Thermal ground truth samples from the DRIVE-IN [13] dataset with side profile.
A.6 Extended Thermalization Ablation Study
@@
Given that the regularizers are solely defined for the synthetic images, we fix $\lambda_{T}=1$ to ensure that the regularizer is on the same scale as the MSE of the real images and vary $\lambda_{W}$ for our ablation. We train all models with the same setup and random seeds. The FID implementation is the default PyTorch-Ignite [16] implementation. Here, we extend the ablation study table in the main document which only displays the best result for $\lambda_{W}=0.01C$. In particular, we display the results of a grid search for $\lambda_{W}$ and $\lambda_{R}$ in Table 6. As described in the main document, we calculate the mean FID and its standard deviation for five different subsets of the T-FAKE dataset. Moreover, we use the same setup to compare the perceptual quality of the ‘cold’ and the ‘warm’ setup of our final T-FAKE dataset, see Table 7. Figure 10 shows T-FAKE samples with the ‘cold’ and ‘warm’ variants. In addition, we present some samples generated with different regularization configurations in Figure 11 for the ‘cold’ setup and in Figure 12 for the ‘warm’ setup. Here, we can visually see the impact of the different regularizers.
@@
$\lambda_{W}=1C$1665 $\pm$ .0030 .3375 $\pm$ .0056
$\lambda_{W}=0.1C$1753 $\pm$ .0018 .3654 $\pm$ .0032
$\lambda_{W}=0.01C$1598 $\pm$ .0041 .3146 $\pm$ .0092
$\lambda_{W}=0$1706 $\pm$ .0029 .5028 $\pm$ .0054
Table 6: Impact of the regularization parameters on the perceptual quality measured with the FID ($\downarrow$) with $C=(5\cdot 8^{2})^{-1}$.
@@
Figure 11: Regularization impact on images for ‘cold’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.01C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.01C$, $\lambda_{R}=1$) (top to bottom).
Figure 12: Regularization impact on images for ‘warm’ setup: No regularization ($\lambda_{W}=0$, $\lambda_{R}=0$), only patch-based ($\lambda_{W}=0.01C$, $\lambda_{R}=0$), only segmentation-based ($\lambda_{W}=0$, $\lambda_{R}=1$), and final model ($\lambda_{W}=0.01C$, $\lambda_{R}=1$) (top to bottom).
Metric Method Training Dataset High Low Side Front Full
NME W/H $\downarrow$ FAKE[61, 62] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
FAKE* FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
Thermal Only FAKE + T-FAKE 0.0832 0.1334 0.0677 0.1503 0.1090
RGB + Thermal FAKE + T-FAKE 0.0740 0.1346 0.0684 0.1420 0.1051
Table 8: Ablation results on CHARLOTTE dataset splits. Pre-processing with the pre-processing stack for RGB landmarkers is indicated by *. The confidence threshold has been set to infinity. RGB + Thermal and Thermal Only models have been finetuned from the FAKE models.
@@ B.2 Landmarking Implementation Details
For training, we include random landmark positions on a texture dataset [17] as negative examples into the dataset to increase the learned uncertainty $\sigma^{2}$ on images without faces. On thermal images, we fill in the background with random textures from the texture dataset [17] with a probability of 0.25. During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes and face scales. We downsample iteratively with a factor of $0.75$ until the height or the width reaches $224$. For each image scale, we run our model on sliding windows of size $224\times 224$ with a stride of $20$. For our final landmark prediction, we pool all predictions and use the landmark with the smallest predicted standard deviation across all scales and all sliding windows. For training the landmarker $T_{\psi}$, we finetune a model that has been pre-trained on the original FAKE dataset for the sparse landmarker for the results in Table 8. During refinement, thermal images are used with a probability of 0.4 split with equal probability for ‘cold’ and ‘warm’ conditions. We finetune for 100 epochs with a learning rate of 0.0004, Adam optimizer with weight decay, batchsize of 512 and OneCycleLR scheduler. Our final model is trained for 4000 epochs, a learning rate of 0.001 on FAKE and T-FAKE (p=0.4).
Augmentations details. We use of spatial augmentations, including random shear, rotations, resizing, and cropping to allow the landmarker to learn a large variation of face orientations without the need for a dedicated face detector. Specifically, we use geometric augmentations which apply random rotations (up to ±45°) and random shear (±7°) to both the image and landmarks, preserving geometric consistency. In addition, a random resized cropping operation is performed, with the cropped region size scaling between $40\%$ and $200\%$ of the original image and an aspect ratio ranging between $\frac{3}{4}$ and $\frac{4}{3}$. Furthermore, we use photometric transformations and random Gaussian smoothing. The thermal images are randomly jittered to simulate temperature variations and thermal images are randomly inverted with a probability of 0.1. Additionally, random noise is applied with a probability of 0.2 to simulate sensor noise. See Figure 18 for image examples with augmentations applied.
@@ B.3 Label Adaptation Implementation Details
Label Adaptation was trained for each method on the predictions on all detected faces on a random 1000 image CHARLOTTE split. We train a model $T_{\zeta}$ for 2000 epochs with a learning rate 0.002, OneCycleLR and Adam optimizer. As landmark augmentations, we apply random rotation up to 45° as well as random shearing during training. The label adaptation network is a five-layer perceptron with fully connected layers that takes the predicted landmarks together with the resize factor as input. The latter accounts for varying degrees of quantization at different image sizes in the CHARLOTTE ground truth. The label adaptation generally handles even outlier predictions but can also contain fail cases, (see Figure 15).
@@
RGB Model Inference. We use two different pre-processing approaches to include landmarkers solely developed for RGB images into the evaluation. Firstly, gray-value images, where the temperature between 20° and 45° is normalized and, secondly the pre-processing stack proposed in [13]. The pre-processing stack consists of temperature clamping between 20°C and 45°C, unsharp masking with two sets of parameters with and without temperature inversion. The reported landmarks are the averages over all detected faces. This simple pre-processing stack is a simple method for boosting RGB landmarker performance for thermal images [13]. As a result, we can include RGB landmarkers as a baseline for thermal landmarking models.
B.4 Landmarking Ablation Study
To study the impact of our thermal data, we report the CHARLOTTE results of our landmarker trained with i) RGB images only, ii) finetuned with thermal images and iii) finetuned with both FAKE and T-FAKE. Again, we use label adaptation for all variations. Table 8 shows the results. Training with the T-FAKE dataset significantly improves the accuracy of thermal landmarking across all conditions. In addition, multimodal training with the FAKE and T-FAKE datasets leads to better thermal landmarking performance than finetuning only with the T-FAKE dataset.
B.5 Inference Ablation
@@
We analyze the impact of our inference strategy, see Table 9. We compare inference on i.) the complete image scaled to $224\times 224$ (whole image), ii.) followed by refinement on a bounding box computed from the predictions obtained with i.) and finally iii.) with the sliding window approach described in the main document. Method iii.) produces the best results over all images except for Charlotte low, where ii.) performs slightly better while at the same time being also suitable for real time estimation.
@@
Figure 14: Results on examples from the CHARLOTTE dataset [3] with different RGB and thermal predictors and our models. For images without landmarks, no faces were detected. The performance of RGB methods can be greatly improved when the images are inverted or sharpened indicated by *. The first column shows the limitations in the CHARLOTTE ground truth: profile annotation convention for frontal views (1st row), quantization artefacts for low resolution images (4th row), translated annotations (last row).
Method Inference Time (ms) High Low Side Front Full
Sliding window ($\bar{\sigma}<\infty$) 88.73 0.0740 0.1346 0.0684 0.1420 0.1051
Refined bounding box ($\bar{\sigma}<\infty$) 10.41 0.0847 0.1329 0.0696 0.1494 0.1095
Whole image ($\bar{\sigma}<\infty$) 5.92 0.1091 0.1868 0.0842 0.2139 0.1490
@@
Table 9: NME (W/H) ($\downarrow$) for CHARLOTTE splits with different strategies for landmark computation. The final results are estimated with sliding windows similar to [62], however, we achieved comparable results when computing the landmarks on input images rescaled to $224\times 224$. Here, we do not exclude high-uncertainty images and evaluate all images without failure rate, i.e., $\bar{\sigma}<\infty$. Average inference time per frame on the Full split has been benchmarked on a single NVIDIA H100 80GB GPU with a batch size of 1.
@@ Appendix C Large-Scale Visualization
For a large number of T-FAKE samples, we refer to Figures 16 and 17. Here, we simply use the first 128 images based on the numerical naming convention of the original FAKE dataset.